# A Clinical Tuning Framework for Continuous Kinematic and Impedance Control of a Powered Knee-Ankle Prosthesis

### v1

@@ Abstract
Objective: Configuring a prosthetic leg is an integral part of the fitting process, but the personalization of a multi-modal powered knee-ankle prosthesis is often too complex to realize in a clinical environment. This paper develops both the technical means to individualize a hybrid kinematic-impedance controller for variable-incline walking and sit-stand transitions, and an intuitive Clinical Tuning Interface (CTI) that allows prosthetists to directly modify the controller behavior.
@@
Methods: Utilizing an established method for predicting kinematic gait individuality alongside a new parallel approach for kinetic individuality, we applied tuned characteristics exclusively from level-ground walking to personalize continuous-phase/task models of joint kinematics and impedance. To take advantage of this method, we developed a CTI that translates common clinical tuning parameters into model adjustments. We then conducted a case study involving an above-knee amputee participant where a prosthetist iteratively tuned the prosthesis in a simulated clinical session involving walking and sit-stand transitions.
Results: The prosthetist fully tuned the multi-activity prosthesis controller in under 20 min. Each iteration of tuning (i.e., observation, parameter adjustment, and model reprocessing) took 2 min on average for walking and 1 min on average for sit-stand. The tuned behavior changes were appropriately manifested in the commanded prosthesis torques, both at the tuned tasks and across untuned tasks (inclines).
Conclusion: The CTI leveraged able-bodied trends to efficiently personalize a wide array of walking tasks and sit-stand transitions. A case-study validated the CTI tuning method and demonstrated the efficiency necessary for powered knee-ankle prostheses to become clinically viable.
I Introduction
@@
Powered prosthetic legs have the potential to restore normative, symmetrical motion to lower-limb amputees, but they are not very common, in part because the high technical burden placed on the clinical team [1, 2, 3]. The clinical training of a prosthetist is multi-faceted, covering socket fabrication, component selection, physiology, gait analysis, and prosthesis personalization, but it is notably not an engineering degree [4, 5]. Conversely, the researchers designing and controlling emergent powered prostheses can create devices that mimic the dynamics of an able-bodied leg, but often lack clinical background to meaningfully implement those devices. Much of this gap between development and implementation relates to the complexity of individualizing the motion for a specific user. This study therefore develops clinically accessible tools and methods to individualize robotic prosthetic legs.
@@
Microprocessor-controlled, variable-damping knee prostheses (e.g., the C-Leg from Ottobock) are personalized by adjusting or ‘tuning’ the damping behavior of the knee with clinically-relevant, digital parameters, such as ‘Swing Flexion Angle’ for the knee and ‘Pushoff Torque’ at the ankle [4, 6]. With powered prosthetic legs, however, one or two joints are replaced with motors capable of more versatile behavior, which adds complexity to the individualization. The controllers that generate commands for these motors must generate different behaviors for different ambulation activities [7, 8] and often have many tunable parameters. Configuring these complex controllers for an individual’s needs is often time-intensive and prohibitively difficult for clinicians [6, 9, 10].
For context, powered prosthetic legs often use impedance controllers to emulate biological joints by changing the stiffness, damping, and equilibrium parameters at specific points in the gait cycle according to a finite-state machine (FSM) [11, 9]. Simple spring-like controllers parameterize each state by defining a torque, $\tau$, as a function of joint angle, $\theta$, and angular velocity, $\dot{\theta}$:
@@
Efforts to address tuning complexity can be categorized into three main approaches: the development of tools to assist clinicians in tuning complex controllers and devices, the development of methods for real-time user tuning, and automated tuning methods that do not incorporate direct user or clinical feedback. On a parallel path, researchers are investigating the specificity and utility of preference in assistive device control and in the tuning of powered prosthetic legs, which is a major component of user and clinical feedback [12, 13, 14, 5]. Simon et al. [15] investigated the first—heuristically tuning a powered knee-ankle prosthesis by determining synergies within the 140 different parameters of a multimodal FSM—but the tuning process still took $\sim$5 hours to complete. Welker et al. allowed the user to directly tune their own impedance parameters to reduce the metabolic cost of walking, with mixed success [16]. Lastly, human-in-the-loop (HITL) reinforcement learning methods [17, 12, 18, 19] algorithmically tune impedance parameters based on aspects of the user’s gait while walking on a treadmill. Though the results of HITL tuning are overall encouraging, these studies focus on tuning only one joint during level-ground walking and do not approach the complexity of coordinated joint motion over many tasks. Further, HITL tuning does not have an avenue to incorporate clinical expertise.
@@
To avoid the complexity of tuning a multimodal FSM, some groups are throwing out the FSM paradigm altogether and using convex optimization to solve for continuous impedance parameter trajectories that parameterize the gait cycle [20, 21, 22, 23, 18]. Best et al. developed a hybrid controller that uses variable impedance during stance and kinematic objectives during swing [20] to accurately emulate normative able-bodied kinematics and kinetics during walking over different inclines and speeds. This continuous paradigm used a parametric model to define the joint impedance, but the parameters that define the model are not necessarily interpretable or tunable.
@@
Our previous analysis of able-bodied datasets identified consistent trends in individuality across variations of walking [7] and stair climbing [24], presenting an interesting avenue to personalize a wide number of tasks given a few clinically accessible observations. Similar underlying trends have been validated by other groups [25, 23]. This method isolates joint-level kinematic individuality (e.g., the difference from the population mean) at level ground and uses that baseline individuality to predict individual kinematics at other speeds and inclines, allowing efficient tuning of continuous-task kinematic controllers like [26]. To extend this concept to joint impedance controllers, we generalized kinetic individuality during walking following the protocol from [24], which can be found in Appendix IV. Given this validation, we then develop a method using baseline individuality to efficiently personalize both the kinematic and impedance controllers defined in [20, 22].
@@
The paper is organized as follows. In Section II, we describe how individualization methods can be applied to the control models, introduce the CTI and its tunable parameters, and describe the experimental protocol. In Section III, we present the results of the case study with an above-knee amputee participant, assessing the efficacy of the tuning method to individualize the controller. Section IV concludes with a discussion of the interactions with our clinical partner and their perception of the tuning process.
II Methods
II-A Quantifying and Distributing Kinetic Individuality Across Tasks
To facilitate efficient individualization of joint torques in a clinical setting, we implemented a strategy to quantify and distribute an individual’s kinetic contribution (IKC) across tasks, using only individual characteristics from one task to individualize the entire task space. We performed an offline validation of this prediction method using the kinematic analysis methods of [24] in Appendix IV. This validation utilizes the dataset in [27, 28] which reports lower-limb kinematics and kinetics of ten able-bodied participants walking at multiple inclines (±0°, 5°, and 10°) and speeds (0.8, 1, 1.2 m/s), amongst other activities.
@@ II-A Quantifying and Distributing Kinetic Individuality Across Tasks
We define the IKC for a given joint and task as the difference between the individual’s joint data ($\mathbf{d}$) and the leave-one-out average (LOO; $\bar{\mathbf{d}}$) of the other nine subjects. In mathematical terms, the IKC is calculated by
@@
for gait phase $\varphi=1,\ldots,150$, task $\chi=1,\ldots,N$, and subject $\eta={1,\ldots,10}$, where $N$ is the number of task variations in the dataset. Figure 1 shows the IKC trajectories for one subject.
The objective is to individualize variations of walking (e.g., different speeds/inclines) using a minimal amount of subject-specific data. Therefore, we only use individuality from level ground movement (i.e., baseline, $\chi_{B}$) to predictively individualize non-baseline tasks by adding the calculated IKC to the population average. The work presented in Appendix IV validates the assumption that the baseline (IKC calculated at $\chi_{B}$) provides a good estimate of individuality across all walking tasks.
II-B Walking Controller
@@
For this study, we individualized the data-driven continuous-phase/task controller developed by Best et al. [20]. This hybrid controller utilizes impedance control during stance and kinematic control during swing. The stance impedance controller is defined by functional representations of impedance parameters $K$, $B$, and $\theta_{eq}$. In swing, the angular position references of the knee and ankle are determined by the continuous kinematic model developed by Embry et al. [29]. These piecewise controllers are both parameterized as continuous functions of gait phase $s$ and task $\chi=(\nu,\gamma)$, where $\nu$ is the subject’s walking speed and $\gamma$ is the ground slope. Both controllers are trained on data from Section II-A[27]. We present simplified explanations of each controller here; further details can be found in the respective papers.
II-B1 Stance—Impedance
@@ II-B2 Swing—Kinematic
II-C Sit-Stand Controller
The sit-stand controller follows the same optimization paradigm as the walking controller (Section II-B1), where the continuous stiffness, equilibrium angle, and damping parameters are solved from able-bodied joint torques using fourth-order polynomials. This method, originally developed by Welker et al. [22], uses the global thigh angle to parameterize phase from sitting ($s=0$) to standing ($s=1$), and from 1 back to 0 for stand-to-sit. While this parameterization neatly organizes phase, the differences in normative dynamics between sitting and standing motions makes it difficult for the user to initiate stand-to-sit in practice. This is resolved by decoupling the sit-to-stand and stand-to-sit equilibrium angles at the knee, adding a dependency on joint angle velocity. This can be seen in the partial reformulation of (3):
@@ II-C Sit-Stand Controller
II-D Clinical Tuning Interface
The CTI implements tuning methods that require minimal training and clinical equipment through the careful implementation of meaningful tuning parameters. These parameters are similar to those of common commercial devices, and automatically translate to changes in the existing model parameters of the controller. Because the controllers are designed to provide able-bodied torques to the user, the tunable parameters are expressed as differences from normative values in either percentage or degrees (e.g., 50% more push off torque at the ankle). The allowed range for each parameter was chosen heuristically, with the bounds set at a safe but extreme value. We now discuss the implementation of each tunable parameter in the context of the model re-optimization.
Refer to caption Refer to caption
Figure 2: The left figure is the CTI with adjustable parameters. A dropdown menu allows the clinician to quickly select preset profiles. Saving and loading of a profile can be completed instantaneously, and this functionality was used to switch between tuned and untuned profiles for testing. Generate Model generates only the models with parameter changes to process the models quickly. VAF gauges for each joint allow for immediate knowledge of fit. Export to Model writes the new model to the board on the prosthesis, which can be updated instantly by an engineer when prompted. The right figure is the advance options box, where the clinician can separate out sit-to-stand and stand-to-sit support.
II-D1 Walking - Stance Flexion Resistance
In commercial devices, stance flexion resistance determines how much force it takes to flex the knee at heel strike; this is often changed in passive devices to minimize foot slap [4]. In this powered device, foot slap is not an issue because we actively control the ankle, so the parameter effectively changes how rigid the knee is as the user progresses through weight acceptance in double support. This change is implemented by changing the heel strike constraint (Section II-B1), which is bounded to [-50%, 60%] of the published value. Anecdotally, this was often described as how ‘squishy’ the knee felt at heel strike.
II-D2 Walking - Swing Knee Flexion
@@
II-D3 Walking - Push-Off Torque
@@
II-D4 Sit-to-Stand - Knee Torque
Tuning knee torque during sit-stand transitions is achieved by scaling the reference torque profiles input into the optimization problem. The scaling remains constant throughout each motion; however, sit-to-stand and stand-to-sit motions were tuned independently. Due to the decoupling of the training parameters during tuning and the equilibrium angles during optimization, there were large changes in equilibrium angles while the stiffness and damping values remained relatively stable across different parameter selections. With split tuning, small variations in thigh velocity can lead to abrupt switching between controllers during the stand-to-sit transition, particularly when the standing and sitting parameters differ substantially. To fix this, we linearly interpolate the scaling factor during the last 10% of standing and the first 10% of sitting. This fix is also applied during sitting when the separation between the sit-to-stand and stand-to-sit parameters is greater than 60%, consequently, we constrained the separation to 60%.
II-E Experiment
The experimental protocol is approved by the Institutional Review Board of the University of Michigan (HUM00166976).
II-E1 Subject Preparation
@@
II-E2 Tuning Protocol
The tuning session was designed to mimic a typical session in a clinic, and began with the user walking on a 7 m handrail walkway with the baseline controller until comfortable not using the handrails. Then, the clinician cycled through a series of preassigned profiles which showcased a high and low setting for each parameter, assessing TF01’s preference. These baselines were designed to apply values at ±$\pm 80\%$ of the parameter space (e.g., a parameter that allows up to a ±$\pm 50\%$ change has a HIGH preset of $+40\%$) to calibrate both the user’s and prosthetist’s mental models.
Then, with the initial preferences in mind, the prosthetist iteratively tuned the prosthesis. This communication was similar to an optometrist visit, where the subject was asked to choose between two options until both the user and clinician agreed that the gait was comfortable. The prosthetist tuned the walking controller, then the sit-stand controller.
II-E3 Experimental Protocol
This protocol was run on both the tuned and un-tuned profiles to determine preference. The order of tested profiles was blinded and randomized, and after each test, the subject was asked to choose their preferred condition. The user performs three trials of a 5-time sit-to-stand test (5xSTS), for 15 total sit-stand repetitions. Then, the subject walked on a treadmill for 30 seconds at -10, 0, and 10, where the inclines tested how well the level-ground tuning parameters extrapolated to untuned tasks.
III Results

### v2

@@ Abstract
Objective: Configuring a prosthetic leg is an integral part of the fitting process, but the personalization of a multi-modal powered knee-ankle prosthesis is often too complex to realize in a clinical environment. This paper develops both the technical means to individualize a hybrid kinematic-impedance controller for variable-incline walking and sit-stand transitions, and an intuitive Clinical Tuning Interface (CTI) that allows prosthetists to directly modify the controller behavior. Methods and procedures: Utilizing an established method for predicting kinematic gait individuality alongside a new parallel approach for kinetic individuality, we personalize continuous-phase/task models of joint impedance (during stance) and kinematics (during swing) using tuned characteristics exclusively from level-ground walking. To take advantage of this method, we developed a CTI that translates common clinical tuning parameters into model adjustments for the walking and sit-stand controllers. We then conducted a case study where a prosthetist iteratively tuned the powered prosthesis to an above-knee amputee participant in a simulated clinical session involving sit-stand transitions and level walking, from which incline/decline walking features were automatically calibrated. Results: The prosthetist fully tuned the multi-activity prosthesis controller in under 20 min. Each iteration of tuning (i.e., observation, parameter adjustment, and model reprocessing) took 2 min on average for walking and 1 min on average for sit-stand. The tuned behavior changes were appropriately manifested in the commanded prosthesis torques, both at the manually tuned tasks and automatically tuned tasks (inclines). Conclusion: The CTI leveraged able-bodied trends to efficiently personalize a wide array of walking tasks and sit-stand transitions, demonstrating the efficiency necessary for powered knee-ankle prostheses to become clinically viable.
Clinical impact: This paper introduces a clinical tuning interface that simplifies the tuning process for multimodal robotic prosthetic legs, reducing the time required from several hours to just 20 minutes thus improving clinical feasibility.
{IEEEkeywords}
Prosthetics, Clinical Tuning, Individualized Care, Robotics, Case Study, Translational Engineering
11footnotetext: 1Department of Robotics at the University of Michigan, Ann Arbor, MI 48109, USA. Corresponding email: rdgregg@umich.edu.
[…]
1 Introduction
[…]
@@ 1 Introduction
Powered prosthetic legs have the potential to restore normative, symmetrical motion to lower-limb amputees, but they are not very common, in part because of the high technical burden placed on the clinical team [1, 2, 3]. The clinical training of a prosthetist is multi-faceted, covering socket fabrication, component selection, physiology, gait analysis, and prosthesis personalization, but it is notably not an engineering degree [4]. Conversely, the researchers designing emergent powered prostheses have expertise in biomechanics, robotics, and control, but often lack the clinical background to meaningfully configure those devices to patients. Much of this gap between development and implementation relates to the complexity of individualizing the motion for a specific user. This study therefore develops clinically accessible tools and methods to individualize robotic prosthetic legs.
@@
Microprocessor-controlled, variable-damping prosthetic knees (e.g., the C-Leg from Ottobock) are personalized by adjusting or ‘tuning’ the damping behavior of the knee with clinically-relevant, digital parameters, such as ‘Swing Flexion Angle’. With powered prosthetic legs, however, one or two joints are replaced with motors capable of more versatile behavior, which adds complexity to the individualization. The controllers that generate commands for these motors must generate different behaviors for different ambulation activities [5, 6] and often have many tunable parameters. Configuring these complex controllers for an individual’s needs is often time-intensive and prohibitively difficult for clinicians [7, 8, 9]. For context, powered prosthetic legs often use impedance controllers to emulate biological joints by changing the stiffness, damping, and equilibrium parameters at specific points in the gait cycle according to a finite-state machine (FSM) [10, 8]. Simple spring-like controllers parameterize each state by defining a torque, $\tau$, as a function of joint angle, $\theta$, and angular velocity, $\dot{\theta}$:
@@
Figure 1: Flowchart of the clinical tuning framework presented in this article. A prior and new analysis of able-bodied data validate the paradigm of individual kinematic/kinetic contributions (IKC) to gait. Following this validation, configurable variables are introduced to two existing hybrid kinematic-impedance controllers—one for variable-incline walking and one for sit-to-stand. A novel clinical tuning interface (CTI) is then introduced to enable manual tuning of the level-ground walking and sit-stand activities, followed by automatic tuning of incline/decline walking using the level-ground controller’s kinematic and kinetic IKC. Finally, both tuned and untuned controllers are assessed and compared during sit-stand and walking activities, where the tuned incline controllers were automatically extrapolated from the manually-tuned level-ground controller.
@@
Efforts to address tuning complexity can be categorized into three main approaches: the development of tools to assist clinicians in tuning complex controllers and devices, the development of methods for real-time user tuning, and automated tuning methods that do not incorporate direct user or clinical feedback. On a parallel path, researchers are investigating the specificity and utility of preference in assistive device control and in the tuning of powered prosthetic legs, which is a major component of user and clinical feedback [11, 12, 13, 14]. Simon et al. [15] heuristically tuned a powered knee-ankle prosthesis by determining synergies within the 140 different parameters of a multimodal FSM, but the tuning process still took $\sim$5 hours to complete. Welker et al. allowed the user to directly tune their own impedance parameters to reduce the metabolic cost of walking, with mixed success [16]. Lastly, human-in-the-loop (HITL) reinforcement learning methods [17, 11, 18, 19, 20] algorithmically tune impedance parameters based on aspects of the user’s gait while walking on a treadmill. While HITL results are overall encouraging, they do not allow prosthetist or patient input to the tuning process. Further, these studies focus on tuning only one joint during level-ground walking and do not approach the complexity of coordinated joint motion over many tasks. Further, HITL tuning does not have an avenue to incorporate clinical expertise.
@@
To avoid the complexity of tuning a multimodal FSM, some groups are moving away from the FSM paradigm and using convex optimization to solve for continuous impedance parameter trajectories that parameterize the gait cycle [21, 22, 23, 24, 18]. Best et al. developed a hybrid controller that uses variable impedance during stance and kinematic objectives during swing [21] to accurately emulate normative able-bodied kinematics and kinetics during walking over different inclines and speeds. This continuous paradigm used a parametric model to define the joint impedance, but the parameters that define the model are not necessarily interpretable or tunable.
@@
Our previous analysis of able-bodied datasets identified consistent trends in individuality across variations of walking [5] and stair climbing [25], presenting an interesting avenue to personalize a wide number of tasks given a few clinically accessible observations. Similar underlying trends have been validated by other groups [26, 24]. This method isolates joint-level kinematic individuality (e.g., the difference from the population mean) at level ground and uses that baseline individuality to predict individual kinematics at other speeds and inclines, allowing efficient tuning of continuous-task kinematic controllers like [27]. To extend this concept to joint impedance controllers, we generalized kinetic individuality during walking following the kinematic individuality protocol from [25]. Given this validation, we then develop a method using baseline individuality to efficiently personalize both the kinematic and impedance controllers defined in [21, 23].
@@
The paper is organized as follows. In Section 2, we describe how individualization methods can be applied to the control models, introduce the CTI and its tunable parameters, and describe the experimental protocol (Figure 1). In Section 3, we present the results of the case study with an above-knee amputee participant, assessing the efficacy of the tuning method to individualize the controller. Section 4 concludes with a discussion of the interactions with our clinical partner and their perception of the tuning process.
2 Methodology
Quantifying and Distributing Individuality Across Tasks
To facilitate efficient individualization of joint torques in a clinical setting, we implemented a strategy to quantify and distribute an individual’s kinematic/kinetic contribution (IKC) across tasks, using only individual characteristics from one task (i.e., one discrete speed and incline combination) to individualize the entire range of tasks. We performed an offline validation of kinetic prediction using the kinematic analysis methods of [25] in Appendix Acknowledgment. This validation utilizes the dataset in [28, 29] which reports lower-limb kinematics and kinetics of ten able-bodied participants walking at multiple inclines (±0°, 5°, and 10°) and speeds (0.8, 1, 1.2 m/s), amongst other activities.
@@ 2 Methodology
We define the kinematic or kinetic IKC for a given joint and task as the difference between the individual’s joint data ($\mathbf{d}$) and the leave-one-out average (LOO; $\bar{\mathbf{d}}$) of the other nine subjects. In mathematical terms, the IKC is calculated by
@@
for gait phase $\varphi=1,\ldots,150$, task $\chi=1,\ldots,N$, and subject $\eta={1,\ldots,10}$, where $N$ is the number of task variations in the dataset. Figure 2 shows the kinetic IKC trajectories for one subject.
The objective of our approach is to individualize variations of walking (e.g., different speeds/inclines) using a minimal amount of subject-specific data. Therefore, we only use individuality from level ground movement (i.e., baseline, $\chi_{B}$) to predictively individualize non-baseline tasks (specifically inclines and declines) by adding the calculated IKC (kinetic during stance and kinematic during swing) to the population average (see flowchart in Figure 1). Appendix Acknowledgment shows that the baseline kinetic individuality (IKC calculated at level-ground task $\chi_{B}$) provides a good estimate of kinetic individuality across other walking tasks.
Walking Controller
@@
For this study, we individualized the data-driven continuous-phase/task controller developed by Best et al. [21]. This hybrid controller utilizes impedance control during stance and kinematic control during swing. The stance impedance controller is defined by functional representations of impedance parameters $K$, $B$, and $\theta_{eq}$. In swing, the angular position references of the knee and ankle are determined by the continuous kinematic model developed by Embry et al. [30]. These piecewise controllers are both parameterized as continuous functions of gait phase $s$ and task $\chi=(\nu,\gamma)$, where $\nu$ is the subject’s walking speed and $\gamma$ is the ground slope. Both controllers are trained on data from Section 2[28]. We present simplified explanations of each controller here; further details can be found in the respective papers.
Stance—Impedance
@@
Sit-Stand Controller
The sit-stand controller follows the same optimization paradigm as the walking controller (Section 2), where the continuous stiffness, equilibrium angle, and damping parameters are solved from able-bodied joint torques using fourth-order polynomials. This method, originally developed by Welker et al. [23], uses the global thigh angle to parameterize phase from sitting ($s=0$) to standing ($s=1$), and from 1 back to 0 for stand-to-sit. While this parameterization neatly organizes phase, the differences in normative dynamics between sitting and standing motions makes it difficult for the user to initiate stand-to-sit in practice. This is resolved by decoupling the sit-to-stand and stand-to-sit equilibrium angles at the knee, adding a dependency on joint angle velocity. This can be seen in the partial reformulation of (3):
@@
Clinical Tuning Interface
The CTI (Figure 3) implements tuning methods that require minimal training and clinical equipment through the careful implementation of meaningful tuning parameters for level-ground walking and sit-stand. These parameters are similar to those of common commercial devices, and automatically translate to changes in the existing model parameters of the controller. Because the controllers are designed to provide able-bodied torques to the user, the tunable parameters are expressed as differences from normative values in either percentage or degrees (e.g., 50% more push off torque at the ankle as seen in Figure 3a). The allowed range for each parameter was chosen heuristically, with the bounds set at a safe but extreme value. Figure 3 also shows several usability features, such as preset profiles, easy saving and loading, and gauges that display Variance Accounted For (VAF), a normalized metric of model fit that is similar to the model’s $R^{2}$. We now discuss the implementation of each tunable parameter in the context of the model re-optimization. These tunable parameters define the user’s kinematic or kinetic IKC, represented by $\psi$, which is automatically applied to the incline and decline walking controllers as previously described.
Figure 3: (a) The CTI with adjustable parameters. A dropdown menu allows the clinician to quickly select preset profiles. Saving and loading of a profile can be completed instantaneously, and this functionality was used to switch between tuned and untuned profiles for testing. Generate Model generates only the models with parameter changes to process the models quickly. Variance Accounted For (VAF) gauges for each joint allow for immediate knowledge of fit. Export to Model writes the new model to the board on the prosthesis, which can be updated instantly by an engineer when prompted. (b) The advanced options dialog box, where the clinician can separate out sit-to-stand and stand-to-sit support.
Walking - Stance Flexion Resistance
In commercial devices, stance flexion resistance determines how much force it takes to flex the knee at heel strike; this is often changed in passive devices to minimize foot slap [4]. In this powered device, foot slap is not an issue because we actively control the ankle, so the parameter effectively changes how rigid the knee is as the user progresses through weight acceptance in double support. This change is implemented by changing the heel strike constraint (Section 2), which is bounded to [-50%, 60%] of the published value. Anecdotally, this was often described as how ‘squishy’ the knee felt at heel strike.
Walking - Swing Knee Flexion
@@
Walking - Push-Off Torque
@@
Sit-to-Stand - Knee Torque
Tuning knee torque during sit-stand transitions is achieved by scaling the reference torque profiles input into the optimization problem. The scaling remains constant throughout each motion; however, sit-to-stand and stand-to-sit motions were tuned independently (Figure 3b). Due to the decoupling of the training parameters during tuning and the equilibrium angles during optimization, there were large changes in equilibrium angles while the stiffness and damping values remained relatively stable across different parameter selections. With split tuning, small variations in thigh velocity can lead to abrupt switching between controllers during the stand-to-sit transition, particularly when the standing and sitting parameters differ substantially. To fix this, we linearly interpolate the scaling factor during the last 10% of standing and the first 10% of sitting. This fix is also applied during sitting when the separation between the sit-to-stand and stand-to-sit parameters is greater than 60%, consequently, we constrained the separation to 60%.
Experimental Methods
The experimental protocol was approved by the Institutional Review Board of the University of Michigan (HUM00166976).
Subject Preparation
@@
Tuning Protocol
The tuning session was designed to mimic a typical session in a clinic, and began with the user walking on a level, 7 m handrail walkway with the baseline controller until comfortable not using the handrails. Then, the clinician cycled through a series of preassigned profiles which showcased a high and low setting for each parameter, assessing TF01’s preference. These baselines were designed to apply values at ±$\pm 80\%$ of the parameter space (e.g., a parameter that allows up to a ±$\pm 50\%$ change has a HIGH preset of $+40\%$) to calibrate both the user’s and prosthetist’s mental models.
Then, with the initial preferences in mind, the prosthetist iteratively tuned the prosthesis. This communication was similar to an optometrist visit, where the subject was asked to choose between two options until both the user and clinician agreed that the gait was comfortable. The prosthetist tuned the walking controller, then the sit-stand controller. The manually-tuned IKC from the level-ground walking controller was automatically applied the incline and decline walking controllers.
Assessment Protocol
This protocol was run on both the tuned and un-tuned profiles to determine preference. The order of tested profiles was blinded and randomized, and after each test, the subject was asked to choose their preferred condition. The user performs three trials of a 5-time sit-to-stand test (5xSTS), for 15 total sit-stand repetitions. Then, the subject walked on an instrumented treadmill (Bertec, Columbus OH) for 30 seconds at -10, 0, and 10, where the incline conditions tested the automatically tuned controllers based on level-ground tuning parameters (IKC). While overground walking trials occurred, they were omitted from these results because they did not contribute a large enough number of steady-state strides at consistent speeds.
3 Results